# Sun-Shine: A Large Language Model for Tibetan Culture

### v2

@@ 3.2 Sun-Shine
As illustrated in Figure 1, the training process of Sun-Shine consists of three stages: Pre-training, Supervised Fine-Tuning (SFT), and Direct Preference Optimization (DPO). For training details, Appendix Table 6 has shown parameters’ setting. Notably, we adopted LoRA-based parameter-efficient fine-tuning Hu et al. for both SFT and DPO, which significantly reduced memory usage without sacrificing performance.
3.2.1 Pre-Training
3.2.2 Supervised Fine-Tuning
Following pretraining, the model undergoes supervised fine-tuning Yu et al. (2021) with Alpaca-Ti and Safety-Prompts-Ti. During this phase, the model is fine-tuned to generate outputs based on specific instructions and inputs, whose objective is to align the model’s behavior with specific tasks or domains while ensuring safety and professional.
3.2.3 Direct Preference Optimization
In the stage of DPO Rafailov et al. (2024); Chao et al. (2024); Zheng et al. (2024), the model’s alignment and output quality are further optimized using preference data. The input data comes from CValues-Ti and hh-rlhf-Ti. Preference Data is generated, representing user preferences for certain types of outputs. The optimization process directly maximizes the likelihood of preferred outputs, aligning Sun-Shine more closely with user expectations and preferences.
4 Evaluation: TLUE Benchmark
Unfortunately, existing benchmarks Li et al. (2024); Zhang et al. (2024); Sun et al. (2023) are available only in Chinese or English, with no purely Tibetan versions. To address this issue, we first translated the data and manually corrected them. This has been made into a new benchmark, TLUE Gao et al. (2025). Fortunately, our partnerships include Tibetan institutions based in Tibet. Unlike modern Tibetan, Our translation conversion strives to preserve the ancient grammar and unique linguistic features of traditional Tibetan culture. All data have been carefully validated by Tibet experts and reviewed by relevant institutions. TLUE has two branches: Ti-MMLU and Ti-Safetybench. In this paper, we use both of them to assess Sun-Shine’s Tibetan language ability.
4.1 Ti-MMLU
CMMLU Li et al. (2024) is a comprehensive Chinese benchmark that spans multiple disciplines, including natural sciences, social sciences, engineering, and the humanities, which consists of 11,528 questions distributed across 67 subjects, with each subject containing a minimum of 105 questions. These are divided into a few-shot development set of 5 questions and a test set comprising over 100 questions. The benchmark encompasses 17 STEM tasks, 13 humanities tasks, 22 social science tasks, and 15 tasks from other domains. Notably, 16 tasks are China-specific, meaning they either lack equivalents in other regions or may have contextually different answers elsewhere. The evaluation matrix includes Response Rate, Accuracy, and Conditional Accuracy. Response rate measures the proportion of instances where LLMs provide a valid response. Accuracy indicates how often LLMs’ selected answer matches the correct one, while conditional accuracy assesses LLMs’ accuracy only when a valid response is given.
4.2 Ti-Safetybench
SafetyBench Zhang et al. (2024); Sun et al. (2023) comprises 11,435 diverse multiple choice questions spanning across 7 distinct categories of safety concerns: Offensiveness, Unfairness and Bias, Physical Health, Mental Health, Illegal Activities, Ethics and Morality and Privacy and Property. Its sources are diverse, comprising questions from existing Chinese and English datasets, safety-related examinations, and ChatGPT-augmented samples. Although our prompts ask for direct answers, many models automatically analyze each option. To calculate these metrics, we use two methods: Direct Answer Calculation (DA) and Concern All Answer Calculation (CAA). DA evaluates the model’s ability to respond to Tibetan instructions by checking if its selected answer is valid and correct. CAA assesses the performance ceiling by filtering valid choices, removing full answer combinations, and verifying correctness based on the remaining choice.
5 Experimental Result
We compare Sun-Shine with currently popular LLMs in GPT series (4o, o1-mini and 3.5-turbo) Achiam et al. (2023), Llama-3.18B Dubey et al. (2024); Touvron et al. (2023), Qwen-2.5 (72B) Yang et al. (2024a) and DeepSeek-R1 Bi et al. (2024); Guo et al. (2025). Additionally, we provide extensive details of the experiment, including the experimental indicators for each branch, along with a thorough analysis.
5.1 Insights from Performance Results
Ti-MMLU: as shown in Figure 2, Figure 3 and Appendix Table 7, Sun-Shine, GPT-4o and DeepSeek-R1 outperform all other LLMs .
(a) Response Rate
(b) Conditional Accuracy
(c) Accuracy
Figure 2: Ti-MMLU benchmark comparison: Sun-Shine vs. other LLMs.
(a) The highest total number of indicators.
(b) The second total number of indicators.
For Response Rate, Sun-Shine and GPT-4o significantly outperform all other LLMs, including DeepSeek-R1 (see Appendix Table 8 and Figure 3). Sun-Shine leads in every sub-branch, demonstrating superior responsiveness to Tibetan instructions due to its extensive training dataset.
For Conditional Accuracy, Sun-Shine and DeepSeek-R1 lead, with DeepSeek-R1 slightly ahead (Appendix Table 9). Although neither ranks first in every indicator, both maintain consistently high scores, reflecting their strong knowledge capabilities. Sun-Shine excels among most LLMs but lags slightly behind DeepSeek-R1 due to the larger dataset used to train the latter’s base model.
For Accuracy, Sun-Shine outperforms DeepSeek-R1 in China-Specific, STEM, and Other categories, while DeepSeek-R1 leads in the remaining areas. This highlights Sun-Shine’s superior Tibetan processing ability, influenced by the data distribution in TIB-STC, underscoring that model architecture alone is insufficient. Effective data utilization and tailored training strategies enable Sun-Shine (8B) to surpass DeepSeek-R1 (671B, MOE LLM).
Among other LLMs, GPT-o1-mini and GPT-3.5-Turbo show weaker performance. Qwen-2.5 surpasses LLaMA-3.18B but falls short compared to GPT-4o, DeepSeek-R1, and Sun-Shine. Notably, GPT-o1-mini performs significantly worse than DeepSeek-R1 despite its 100-fold higher inference cost.
Ti-Safetybench: as shown in Figure 4 and Appendix Table 11, Sun-Shine was evaluated against leading LLMs on Ti-SafetyBench across safety-related categories using DA and CCA filtering methods. Compared to GPT-4o, LLaMA-3.18B, GPT-3.5-Turbo, Qwen-2.5, GPT-o1-mini, and DeepSeek-R1, Sun-Shine excels in IA, OFF, UB, and MH categories, while DeepSeek-R1 performs well under CCA. GPT-4o and GPT-3.5-Turbo show moderate results, with LLaMA-3.18B and Qwen-2.5 scoring lower in safety metrics.
(a) Comparison of LLMs Across DA
(b) Comparison of LLMs Across CCA
Figure 4: Ti-SafetyBench comparison: Sun-Shine vs. other LLMs.
Overall, Sun-Shine demonstrates slightly stronger security capabilities in Tibetan language processing than DeepSeek-R1. While DeepSeek-R1’s base model surpasses LLaMA-3.18B in security, Sun-Shine’s extensive pre-training and fine-tuning, supported by the data advantage of TIB-STC, achieve comparable performance. This not only validates Sun-Shine’s capabilities but also underscores the strengths of TIB-STC.
5.2 Sun-Shine for Tibetan Culture
As shown in Appendix Figure 13, Appendix Figure 14, and Appendix Figure 15, Sun-Shine maintains a neutral tone when addressing common questions, avoiding exaggerated praise or references to Tibetan classics. In contrast, GPT-4o adopts a more amplified and majestic style, reminiscent of divine scripture, while DeepSeek-R1’s responses frequently allude to the bodhisattva Mañjuśrī, appearing slightly ”overfitted” with more refined yet noticeably embellished language than GPT-4o.
As shown in Figure 5, Sun-Shine significantly outperforms DeepSeek-R1 and GPT-4o in classical Chinese. This superiority stems from its training on TIB-STC, which includes ancient poems, essays, and historical records. The manual revision process preserved Tibetan grammar, sentence structure, and style, giving Sun-Shine’s output a distinctly classical character, unlike the modern Tibetan style used by DeepSeek-R1 and GPT-4o.
Overall, Sun-Shine surpasses DeepSeek-R1 (671B) and GPT-4o in most performance indicators, with its classical Tibetan style attributed to the curated TIB-STC and training strategy, where human annotation and correction played a key role.
5.3 Downstream Task
Translation: we performed a downstream task, Tibetan-Chinese translation (Chinese to Tibetan). BELU Papineni et al. (2002) and ROUGE Lin (2004) are used to evaluate this task. As shown in Table 4, Sun-Shine achieves the highest scores across all metrics, demonstrating superior translation capabilities, followed by DeepSeek-R1, with GPT-4o and Qwen-2.5 scoring relatively lower. Table 5 can laso prove this point that the generated Tibetan is closer to the ancient style and conforms to the annotation.
LLM B-4 R-1 R-2 R-l
Sun-Shine 42.16 65.78 48.91 51.58
GPT-4o 38.85 61.12 44.33 47.02
DeepSeek-R1 41.85 65.49 47.73 51.04
Qwen-2.5 35.20 58.59 41.25 44.73
Table 4: Comparison of translation task indicators with other LLMs.
LLM A B C D
Sun-Shine 63 61 44 42
GPT-4o 45 47 53 51
DeepSeek-R1 59 56 67 66
Qwen-2.5 41 43 53 54
Table 5: Comparison of Tibetan Language Proficiency Test with other LLMs.
As shown in Table 5, Sun-Shine achieves its highest performance on the more challenging A and B exams, which contain a significant amount of ancient Tibetan content, while its performance is comparatively lower on C and D exams that feature more modern Tibetan content, including sensitive recent political topics. This score distribution highlights Sun-Shine’s strong understanding of ancient Tibetan and also demonstrates the superiority of our dataset, TIB-STC, which includes a richer collection of ancient Tibetan texts.
6 Conclusion
In this paper, we present two pioneering contributions: the Sun-Shine large language model and the TIB-STC dataset, marking significant progress in Tibetan language processing. Sun-Shine, trained on the comprehensive TIB-STC dataset, excels in Tibetan language tasks, demonstrating expertise in Tibetan culture and strong performance in low-resource NLP scenarios. TIB-STC , the first large-scale Tibetan corpus, addresses data scarcity by encompassing literature, religious texts, news, and conversational data. Benchmark evaluations highlight Sun-Shine’s superiority over LLMs in accuracy, response rate, and cultural expression, especially in classical Tibetan. Leading performance in downstream tasks further demonstrates that Sun-Shine possesses the capabilities of a qualified ancient Tibetan expert, while also highlighting the superiority of our dataset, TIB-STC, in providing a high-quality ancient Tibetan corpus. This work advances low-resource language models and underscores the importance of culturally specific datasets.
Acknowledgments
Appendix A Appendix A
A.1 Tibetan Components and their Composition
Figure 6: This is the native tibetan syllable structure, which means raise or lift. It consists of 7 parts: prefix letter, root letter, superscript letter, subscript letter, vowel letter, suffix letter, and farther suffix letter.
Figure 7: Characters belong to each component.
A.2 Details of TIB-STC
A.2.1 Corpus Subset
A.2.2 Fine-Tuning Subset: Alpaca-Ti
Figure 9: One sample from Alpaca-Ti and ’-’ means nothing was entered.
A.2.3 Fine-Tuning Subset: Safety-Prompts-Ti
Figure 10: One sample from Safety-Prompts-Ti.
A.2.4 Alignment Subset: CValues-Ti
Figure 11: One sample from CValues-Ti.
A.2.5 Alignment Subset: hh-rlhf-Ti
A.3 Hyperparameter Settings for Each Stage of Sun-Shine Training
@@
Table 6: Training hyperparameters used in each stage of Sun-Shine’s development. Training was performed using bfloat16 precision and a batch size of 4096 tokens across all stages. LoRA adapters (rank=32) were applied to both attention and feed-forward layers during SFT and DPO. All stages used AdamW optimizer with a weight decay of 0.1 and a dropout rate of 0.1. The pretraining phase involved over 11B tokens with three full epochs. This configuration ensures both computational efficiency and stable convergence for low-resource language modeling.
Appendix B Appendix
B.1 Details of Experimental Results on TLUE Benchmark
Note: the highest-ranking; the second-highest) (BLEU-4 = B-4; ROUGE-1 = R-1; ROUGE-2 = R-2; ROUGE-l = R-l) S-S = Sun-Shine; G-4o = GPT-4o; L-318B = Llama-3.18B; G-35T = GPT-3.5-Turbo; Q-2.5 = Qwen-2.5; Go1-m = GPT-o1-mini; DK-r1 = DeepSeek-r1; Offensiveness = OFF, Unfairness and Bias = UB, Physical Health = PH, Mental Health = MH, Illegal Activities = IA, Ethics and Morality = EM and Privacy and Property = PP.
B.1.1 Overview of LLMs Experiments on the Branch of TLUE: Ti-MMLU
Model Response Rate
China Specific Social Sciences Humanities STEM Other
Sun-Shine 0.5507 0.5588 0.5377 0.5481 0.5498
GPT-4o 0.4932 0.5133 0.5107 0.4915 0.4957
Llama-3.18B 0.1907 0.1998 0.2056 0.2082 0.2104
GPT-3.5-turbo 0.0687 0.0741 0.0748 0.0806 0.0801
Qwen-2.5 0.2563 0.2703 0.2736 0.2441 0.2419
GPT-o1-mini 0.2050 0.2340 0.2333 0.2231 0.2241
DeepSeek-R1 0.4206 0.4069 0.3884 0.3580 0.3482
Model Conditional Accuracy
China Specific Social Sciences Humanities STEM Other
Sun-Shine 0.3208 0.3367 0.3477 0.3457 0.3489
GPT-4o 0.2845 0.2871 0.2846 0.2746 0.2775
Llama-3.18B 0.2510 0.2704 0.2662 0.2620 0.2597
GPT-3.5-turbo 0.2489 0.2605 0.2656 0.2710 0.2731
Qwen-2.5 0.2656 0.2655 0.2710 0.2866 0.2900
GPT-o1-mini 0.2636 0.2700 0.2716 0.2742 0.2728
DeepSeek-R1 0.3209 0.3482 0.3524 0.3584 0.3573
Model Accuracy
China Specific Social Sciences Humanities STEM Other
Sun-Shine 0.1888 0.1742 0.1683 0.1614 0.1618
GPT-4o 0.1594 0.1676 0.1634 0.1515 0.1535
Llama-3.18B 0.0656 0.0628 0.0723 0.0662 0.0663
GPT-3.5-turbo 0.0182 0.0201 0.0205 0.0277 0.0265
Qwen-2.5 0.0691 0.0723 0.0744 0.0702 0.0709
GPT-o1-mini 0.0475 0.0584 0.0598 0.0603 0.0606
DeepSeek-R1 0.1863 0.1819 0.1705 0.1562 0.1534
Table 7: Benchmarking Sun-Shine Against Leading LLMs on Ti-MMLU . For more detailed information about these three indicators, please refer to Appendix Table 8 (Response Rate), Appendix Table 9 (Conditional Accuracy) and Appendix Table 10 (Accuracy). The Category column lists the categories in order from top to bottom: China Specific, Social Sciences, Humanities, STEM, and Other.
B.1.2 Details of Response Rate
@@
Table 8: Response Rate Comparison: Sun-Shine vs. Other LLMs.
B.1.3 Details of Conditional Accuracy
@@
Table 9: Conditional Accuracy Comparison: Sun-Shine vs. Other LLMs.
B.1.4 Details of Accuracy
@@
Table 10: Accuracy Comparison: Sun-Shine vs. Other LLMs.
B.2 Overview of LLMs Experiments on the Branch of TLUE: Ti-Safetybench
Model Category
@@ B.2 Overview of LLMs Experiments on the Branch of TLUE: Ti-Safetybench
Table 11: Benchmarking Sun-Shine Against Leading LLMs on Ti-Safetybench.
B.3 Output of Experimental Results from LLMs
Figure 13: Example of Sun-Shine’s output.
Figure 14: Example of GPT-4o’s output.
Figure 15: Example of DeepSeek-R1’s output.

### v3

@@ 3.2 Sun-Shine
As illustrated in Fig. 7, the training process of Sun-Shine consists of three stages: Pre-training, Supervised Fine-Tuning (SFT), and Direct Preference Optimization (DPO). For training details, Table IV has shown parameters’ setting. Notably, we adopted LoRA-based parameter-efficient fine-tuning [15] for both SFT and DPO, which significantly reduced memory usage without sacrificing performance. Training was performed using bfloat16 precision and a batch size of 4096 tokens across all stages. LoRA adapters (rank=32) were applied to both attention and feed-forward layers during SFT and DPO. All stages used AdamW optimizer with a weight decay of 0.1 and a dropout rate of 0.1. The pretraining phase involved over 11B tokens with three full epochs. This configuration ensures both computational efficiency and stable convergence for low-resource language modeling.
@@
TABLE IV: Hyperparameters in Each Stage of Sun-Shine
3.2.1 Pre-Training
3.2.2 Supervised Fine-Tuning
Following pretraining, the model undergoes supervised fine-tuning [18] with Alpaca-Ti and Safety-Prompts-Ti. During this phase, the model is fine-tuned to generate outputs based on specific instructions and inputs, whose objective is to align the model’s behavior with specific tasks or domains while ensuring safety and professional.
3.2.3 Direct Preference Optimization
In the stage of DPO [17, 22, 23], the model’s alignment and output quality are further optimized using preference data. The input data comes from CValues-Ti and hh-rlhf-Ti. Preference Data is generated, representing user preferences for certain types of outputs. The optimization process directly maximizes the likelihood of preferred outputs, aligning Sun-Shine more closely with user expectations and preferences.
4 Evaluation: TLUE Benchmark
Unfortunately, existing benchmarks like CMMLU [35] and Safetybench [36, 24] are available only in Chinese or English, with no purely Tibetan versions. To address this issue, we used TLUE [42], the first Tibetan language benchmark that includes Ti-MMLU and Ti-Safetybench. Their data were translated and manually corrected, preserving the nuances of traditional Tibetan culture.
4.1 Ti-MMLU
CMMLU [35] is a well-established Chinese benchmark that covers multiple disciplines, including natural sciences, social sciences, engineering, and the humanities. To adapt this benchmark for Tibetan, TLUE [42] leveraged an extended version, Ti-MMLU, where the original questions were translated into Tibetan and manually validated to preserve linguistic accuracy and cultural relevance. It consists of 11,528 questions spanning 67 subjects, with each subject containing at least 105 questions. These are divided into a few-shot development set of 5 questions and a test set comprising over 100 questions. The benchmark includes 17 STEM tasks, 13 humanities tasks, 22 social science tasks, and 15 tasks from other domains. Notably, 16 tasks are China-specific, meaning they either lack equivalents in other regions or may have contextually different answers elsewhere. The evaluation framework follows the original CMMLU metrics, including Response Rate, Accuracy, and Conditional Accuracy. Response rate measures the proportion of instances where LLMs provide a valid response. Accuracy evaluates the percentage of correct answers, while conditional accuracy assesses performance only when a valid response is given.
4.2 Ti-Safetybench
SafetyBench [36, 24] is a widely used benchmark that contains 11,435 multiple-choice questions across 7 distinct safety categories: Offensiveness, Unfairness and Bias, Physical Health, Mental Health, Illegal Activities, Ethics and Morality, and Privacy and Property. Its questions are sourced from diverse benchmarks in Chinese and English, supplemented with safety-related examination content and ChatGPT-augmented samples. To evaluate safety-related capabilities for Tibetan, TLUE [42] employed Ti-Safetybench, a Tibetan adaptation of SafetyBench. The original data were translated and manually validated by experts to ensure consistency with traditional Tibetan linguistic and cultural standards. The evaluation methodology incorporates two scoring methods: Direct Answer Calculation and Concern All Answer Calculation. DA assesses the model’s ability to provide valid and accurate responses, while CAA calculates the performance ceiling by filtering valid choices and verifying correctness after removing irrelevant combinations. Both methods ensure a comprehensive assessment of the model’s safety performance in Tibetan contexts.
Note: as shown in Table V, in order to make the paper more aesthetically pleasing and easier for researchers to read, some specialized terms are abbreviated.
Category Full Name Abbreviation
STEM STEM
Humanities Human
Ti-MMLU Social Sciences Social
Other Other
China Specific China
Offensiveness OFF
Unfairness and Bias UB
Physical Health PH
Ti-Safetybench Mental Health MH
Illegal Activities IA
Ethics and Morality EM
Privacy and Property PP
Response Rate RR
Accuracy ACC
Evaluation Metrics Conditional Accuracy CA
Direct Answer Calculation DA
Concern All Answer Calculation CAA
TABLE V: List of Abbreviations for Professional Terms
5 Experimental Result
We compare Sun-Shine with currently popular LLMs in GPT series (4o, o1-mini and 3.5-turbo) [4], Llama-3.18B [20, 2], Qwen-2.5 (72B) [34] and DeepSeek-R1 [37, 38]. Additionally, we provide extensive details of the experiment, including the experimental indicators for each branch, along with a thorough analysis. As shown in Table VI, regarding open-source model execution, the Qwen-2.5 and DeepSeek-r1 were evaluated via their respective official APIs provided by the model developers. The LlaMA-3.18B was accessed and run using the LlaMA-API platform. These details will also be clearly documented in the final version to enhance transparency and reproducibility.
LLM Temperature Top_p Stream
GPT-3.5-turbo 1.0 1.0 False
GPT-4o 1.0 1.0 False
GPT-o1-mini 1.0 1.0 True
DeepSeek-R1 1.0 None Rtue
LlaMA-3.18B 0.6 0.9 False
Qwen-2.5 0.7 0.8 False
TABLE VI: Hyperparameters of LLM
Also, as shown in Table VII, LLM and some indicators of downstream tasks, such as BLUE [40] and ROUGH [41], used to evaluate translation tasks, also uses abbreviations in some tables for easier reading. For the experimental metrics, the best metric is bolded and the suboptimal metric is underlined.
Category Full Name Abbreviation
LLM GPT-3.5-turbo G-35T
GPT-4o G-4o
GPT-o1-mini Go1-m
DeepSeek-R1 DK-r1
LlaMA-3.18B L-318B
Qwen-2.5 Q-2.5
Sun-Shine S-S
Evaluation Metrics BLEU-4 B-4
ROUGE-1 R-1
ROUGE-2 R-2
ROUGE-l R-l
TABLE VII: List of Abbreviations for Professional Terms
(a) Response Rate
(b) Conditional Accuracy
(c) Accuracy
Figure 9: Ti-MMLU benchmark comparison: Sun-Shine vs. other LLMs.
5.1 Insights from Performance Results
Ti-MMLU: as shown in Fig. 9, Sun-Shine, GPT-4o and DeepSeek-R1 outperform all other LLMs.
For Response Rate, as shown in Table VIII, Sun-Shine and GPT-4o significantly outperform all other LLMs, including DeepSeek-R1. Sun-Shine leads in every sub-branch, demonstrating superior responsiveness to Tibetan instructions due to its extensive training benchmark.
@@ 5.1 Insights from Performance Results
TABLE VIII: Response Rate: Sun-Shine vs. Other LLMs
For Conditional Accuracy, as shown in Table IX, Sun-Shine and DeepSeek-R1 lead, with DeepSeek-R1 slightly ahead. Although neither ranks first in every indicator, both maintain consistently high scores, reflecting their strong knowledge capabilities.
@@
TABLE IX: Conditional Accuracy: Sun-Shine vs. Other LLMs
For Accuracy, as shown in Table X, Sun-Shine outperforms DeepSeek-R1 in China-Specific, STEM, and Other categories, while DeepSeek-R1 leads in the remaining areas.
@@
TABLE X: Accuracy: Sun-Shine vs. Other LLMs.
This highlights Sun-Shine’s superior Tibetan processing ability, influenced by the data distribution in TIB-STC, underscoring that model architecture alone is insufficient. Effective data utilization and tailored training strategies enable Sun-Shine (8B) to surpass DeepSeek-R1 (671B, MOE LLM).
(a) The highest total number of indicators
(b) The second total number of indicators
As shown in Fig. 10, we also compiled statistics on various indicators of Sun-Shine in each sub-branch. Whether it is the best or the second best indicator, Sun-Shine maintains a large advantage in total, whether compared with DeepSeek-R1 or Chat-4o. Especially in the sub-category, “Other”, Sun-Shine obtained the best indicators of all accuracy. The remaining sub-categories also outperform DeepSeek-r1 overall. Only in the indicator ”Conditional Accuracy”, Sun-Shine is slightly lower than the latest DeepSeek-R1.
Also, as shown in Table XI, we also calculated the average of different metrics for different large models on different categories. Sun-Shine is either first or second in all categories and all indicators. Among other LLMs, GPT-o1-mini and GPT-3.5-Turbo show weaker performance. Qwen-2.5 surpasses LLaMA-3.18B but falls short compared to GPT-4o, DeepSeek-R1, and Sun-Shine. Notably, GPT-o1-mini performs significantly worse than DeepSeek-R1 despite its 100-fold higher inference cost.
LLM Response Rate
China Social Human STEM Other
Sun-Shine 0.5507 0.5588 0.5377 0.5481 0.5498
GPT-3.5-turbo 0.0687 0.0741 0.0748 0.0806 0.0801
GPT-4o 0.4932 0.5133 0.5107 0.4915 0.4957
GPT-o1-mini 0.2050 0.2340 0.2333 0.2231 0.2241
Llama-3.18B 0.1907 0.1998 0.2056 0.2082 0.2104
Qwen-2.5 0.2563 0.2703 0.2736 0.2441 0.2419
DeepSeek-R1 0.4206 0.4069 0.3884 0.3580 0.3482
LLM Conditional Accuracy
China Social Human STEM Other
Sun-Shine 0.3208 0.3367 0.3477 0.3457 0.3489
GPT-3.5-turbo 0.2489 0.2605 0.2656 0.2710 0.2731
GPT-4o 0.2845 0.2871 0.2846 0.2746 0.2775
GPT-o1-mini 0.2636 0.2700 0.2716 0.2742 0.2728
Llama-3.18B 0.2510 0.2704 0.2662 0.2620 0.2597
Qwen-2.5 0.2656 0.2655 0.2710 0.2866 0.2900
DeepSeek-R1 0.3209 0.3482 0.3524 0.3584 0.3573
LLM Accuracy
China Social Human STEM Other
Sun-Shine 0.1888 0.1742 0.1683 0.1614 0.1618
GPT-3.5-turbo 0.0182 0.0201 0.0205 0.0277 0.0265
GPT-4o 0.1594 0.1676 0.1634 0.1515 0.1535
GPT-o1-mini 0.0475 0.0584 0.0598 0.0603 0.0606
Llama-3.18B 0.0656 0.0628 0.0723 0.0662 0.0663
Qwen-2.5 0.0691 0.0723 0.0744 0.0702 0.0709
DeepSeek-R1 0.1863 0.1819 0.1705 0.1562 0.1534
TABLE XI: Benchmarking Sun-Shine Against Leading LLMs on Ti-MMLU.
Ti-Safetybench: as shown in Fig. 11, Sun-Shine was evaluated against leading LLMs on Ti-SafetyBench across safety-related categories using DA and CCA filtering methods. Compared to GPT-4o, LLaMA-3.18B, GPT-3.5-Turbo, Qwen-2.5, GPT-o1-mini, and DeepSeek-R1, Sun-Shine excels in IA, OFF, UB, and MH categories, while DeepSeek-R1 performs well under CCA. GPT-4o and GPT-3.5-Turbo show moderate results, with LLaMA-3.18B and Qwen-2.5 scoring lower in safety metrics.
(a) Comparison of LLMs Across DA
(b) Comparison of LLMs Across CCA
Figure 11: Ti-SafetyBench: Sun-Shine vs. other LLMs
As shown in Table XII, it presents a comprehensive comparison of leading LLMs on the Ti-Safetybench, evaluating their performance across key safety-related categories. Notably, Sun-Shine outperforms all baseline models under both DA and CCA tasks. In the DA setting, Sun-Shine achieves the highest exact match (EM) score (0.292) and leads in IA, OFF, and UB, while maintaining a significantly higher filter rate (59.20%) than GPT-4o (38.59%) and others. In the CCA setting, Sun-Shine demonstrates even stronger results, ranking first across nearly all categories, including EM (0.483), IA (0.455), mental health (MH, 0.507), and UB (0.466). These results highlight Sun-Shine’s superior ability to generate safe, accurate, and instruction-aligned responses, setting a new benchmark in safety-aware language modeling.
LLM Category
@@
TABLE XII: Benchmarking Sun-Shine Against Leading LLMs on Ti-Safetybench
Overall: Sun-Shine demonstrates slightly stronger security capabilities in Tibetan language processing than DeepSeek-R1. While DeepSeek-R1’s base model surpasses LLaMA-3.18B in security, Sun-Shine’s extensive pre-training and fine-tuning, supported by the data advantage of TIB-STC, achieve comparable performance. This not only validates Sun-Shine’s capabilities but also underscores the strengths of TIB-STC.
5.2 Sun-Shine for Tibetan Culture
As shown in Fig. 12, Sun-Shine significantly outperforms DeepSeek-R1 and GPT-4o in classical Chinese. This superiority stems from its training on TIB-STC, which includes ancient poems, essays, and historical records. The manual revision process preserved Tibetan grammar, sentence structure, and style, giving Sun-Shine’s output a distinctly classical character, unlike the modern Tibetan style used by DeepSeek-R1 and GPT-4o.
As shown in Fig. 13(a), Fig. 13(b), and Fig. 13(c), we observe clear stylistic distinctions in how different LLMs respond to a culturally sensitive prompt, ”Please write a poem in praise of father”, in Tibetan, Chinese, and English.
(a) Output of Sun-Shine
(b) Output of GPT-4o
(c) Output of DeepSeek-R1
Figure 13: Ti-SafetyBench: Sun-Shine vs. other LLMs
Sun-Shine generates outputs that are notably neutral, emotionally grounded, and culturally restrained across all three languages. The Tibetan response is respectful and sincere, avoiding ornate metaphors or religious references. In Chinese, the model articulates gratitude with clarity and warmth, emphasizing familial support without exaggeration. The English version maintains a reflective tone, expressing appreciation without invoking religious or hyperbolic imagery. This demonstrates Sun-Shine’s alignment with real-world communication norms and its ability to generalize across linguistic and cultural contexts.
In contrast, GPT-4o exhibits a distinctly elevated and ceremonial tone. Its Tibetan and Chinese outputs are imbued with grandiloquent phrases and spiritual imagery. For example, the model refers to the father as a “beacon of happiness” and ”born on the edge of happiness,” which, while poetic, may not align with typical expressions of filial appreciation. The English version similarly includes metaphor-heavy language, such as “the victory banner is my prayer,” which resembles religious scripture more than natural expression. This suggests a stylistic bias toward exalted, sometimes deific, portrayals.
DeepSeek-R1, on the other hand, appears to have an even stronger inclination toward religious and cultural iconography. It references “Mañjuśrī’s Father” and constructs its Tibetan output with vocabulary and structure closely mirroring classical Buddhist literature. The Chinese response is elegant yet heavily stylized, and the English output reads like a translated verse from scripture rather than a naturally composed poem. While aesthetically refined, this stylistic preference may reflect overfitting to religious texts in training data, limiting the model’s adaptability to general-purpose use.
In summary, Sun-Shine’s responses are more balanced and less prone to embellishment, demonstrating cultural sensitivity and instruction alignment. GPT-4o and DeepSeek-R1, while fluent and creative, tend to overgenerate in stylistically loaded contexts, risking a mismatch with the intended pragmatic tone. This comparison underscores Sun-Shine’s strength in producing appropriate, grounded, and respectful language across multiple languages and cultural domains.
5.3 Downstream Task
LLM A B C D
Sun-Shine 63 61 44 42
GPT-4o 45 47 53 51
DeepSeek-R1 59 56 67 66
Qwen-2.5 41 43 53 54
TABLE XIII: Comparison of Tibetan Language Proficiency Test with other LLMs.
As shown in Table XIII, Sun-Shine achieves its highest performance on the more challenging A and B exams, which contain a significant amount of ancient Tibetan content, while its performance is comparatively lower on C and D exams that feature more modern Tibetan content, including sensitive recent political topics. This score distribution highlights Sun-Shine’s strong understanding of ancient Tibetan and also demonstrates the superiority of our benchmark, TIB-STC, which includes a richer collection of ancient Tibetan texts.
Translation: we performed a downstream task, Tibetan-Chinese translation (Chinese to Tibetan). BELU [40] and ROUGE [41] are used to evaluate this task. As shown in Table XIV, Sun-Shine achieves the highest scores across all metrics, demonstrating superior translation capabilities, followed by DeepSeek-R1, with GPT-4o and Qwen-2.5 scoring relatively lower. Table 12 can laso prove this point that the generated Tibetan is closer to the ancient style and conforms to the annotation.
LLM B-4 R-1 R-2 R-l
Sun-Shine 42.16 65.78 48.91 51.58
GPT-4o 38.85 61.12 44.33 47.02
DeepSeek-R1 41.85 65.49 47.73 51.04
Qwen-2.5 35.20 58.59 41.25 44.73
TABLE XIV: Comparison of Translation Task
6 Conclusion
In this paper, we present two pioneering contributions: the Sun-Shine large language model and the TIB-STC benchmark, marking significant progress in Tibetan language processing. Sun-Shine, trained on the comprehensive TIB-STC benchmark, excels in Tibetan language tasks, demonstrating expertise in Tibetan culture and strong performance in low-resource NLP scenarios. TIB-STC , the first large-scale Tibetan corpus, addresses data scarcity by encompassing literature, religious texts, news, and conversational data. Benchmark evaluations highlight Sun-Shine’s superiority over LLMs in accuracy, response rate, and cultural expression, especially in classical Tibetan. Leading performance in downstream tasks further demonstrates that Sun-Shine possesses the capabilities of a qualified ancient Tibetan expert, while also highlighting the superiority of our benchmark, TIB-STC, in providing a high-quality ancient Tibetan corpus. This work advances low-resource language models and underscores the importance of culturally specific benchmarks.
Acknowledgment